You invested 10,000 in two stocks paying 8% and 10% annual interest. At the end of the year the total interest from these investments was 940. How much was invested at each rate?
Answer by addingup(3677)

Of 10,000, a certain amount pays 8%. The rest of the money pays 10%. In other words, x money pays .08, so we can say .08x And what is left of the 10,000 pays .10. So we can say:
--------------------------
.08x+.10*10,000-x= 940 Multiply on left:
.08x+ 1000-x= 940 Subtract 1000 on both sides and do subtraction on left:
-.02x= -60 Divide both sides by -.02:
x= -60/-.02
x= 3000 This is the amount that pays 8% And 10,000-3,000= 7,000 pays 10%.
Let's see:
3,000*.08= 240
7,000*.10= 700
------------
Total......940